The 3rd angle in an isolsoles triangle is 8 degrees less than twice the measure of the equal angle. Find the measure of one of the equal angle.
Thanks for your help

Answer by ankor@dixie-net.com(22740)

The 3rd angle in an isosoles triangle is 8 degrees less than twice the measure of the equal angle. Find the measure of one of the equal angle.
:
Let x = the measure of one of the equal angles
:
It says,"The 3rd angle is 8 degrees less than twice the measure of the equal angle." therefore
(2x - 8) = 3rd angle
:
The three angles = 180
x + x + (2x-8) = 180
:
4x - 8 = 180
:
4x = 180 + 8
:
x = 188/4
:
x = 47 degrees is the measure of one of the equal angles
;
:
Check solution, 3rd angle = 2(47) - 8 = 86 degrees:
47 + 47 + 86 = 180